OpenStudy (anonymous):

I will give medal and fan for help in return :D A used car dealership buys a car for $2600 and then sells it for $4300. What is the percentage increase?

OpenStudy (anonymous):

lets say that 2600$ = 100% now the percentage will be our X , $4300 = X% now 4300x100 / 2600 = 165.38% I'm not sure but i guess this is the right answer or might be , 165,38% - 100%= 65,38% the net percentage increase

OpenStudy (anonymous):

ill try both :)

OpenStudy (anonymous):

they're both right , its just the question that might be Mysterious :)

OpenStudy (anonymous):

it says they are both wrong answers :(

OpenStudy (anonymous):

@tinybookworm

OpenStudy (anonymous):

I think @Ogzii was right. How many decimal places does it want?

OpenStudy (anonymous):

it says to round to the nearest whole number as needed

OpenStudy (anonymous):

Oh have you tried 65% or 165%?

OpenStudy (anonymous):

no

OpenStudy (anonymous):

which one d you think would be correct because i only have 1 more chance

OpenStudy (anonymous):

@tinybookworm

OpenStudy (anonymous):

I think it is 65

OpenStudy (anonymous):

okit was I have some more

OpenStudy (anonymous):

Ok. I will help you if you want

OpenStudy (anonymous):

tell whether the percent change is an increase or a decrease. find the percent change 

OpenStudy (anonymous):

that would be awesome you and @sangya21 have been super helpful today :)

OpenStudy (anonymous):

any ideas on this one?

OpenStudy (anonymous):

Th ank you. I am sure you can tell if it is an increase or decrease

OpenStudy (anonymous):

is it really supposed to be that simple ?

OpenStudy (anonymous):

To calculate the change, find the subtract the new amount by the original one, then divide it by the original. If it is +, it increases. If it is -, then it decreases.

OpenStudy (anonymous):

it was positive so increase :)

OpenStudy (anonymous):

now i have to find the percent change :/

OpenStudy (anonymous):

would it be the number i found when subtracting them?

OpenStudy (anonymous):

No, it would be the difference divide by the original amount, then time 100

OpenStudy (anonymous):

it is increasing 50 = 100% 70 = 140% rate of increasing is 40%

OpenStudy (anonymous):

Spend a little time on this site: http://www.skillsyouneed.com/num/percent-change.html. I found it quite helpful

OpenStudy (anonymous):

wait i meant this: 70/50 times 100

OpenStudy (anonymous):

It should be \(\frac{70-50}{50}\times 100\). @Ogzii was right. You can also use his way to solve this

OpenStudy (anonymous):

oh ok

OpenStudy (anonymous):

so im going to get 40%

OpenStudy (anonymous):

either way?

OpenStudy (anonymous):

\(\frac {70}{50}\times 100 -100\)

OpenStudy (anonymous):

oh cool

OpenStudy (anonymous):

i am writting my answer down give me 1 min

OpenStudy (anonymous):

Ok

OpenStudy (anonymous):

uhmm i have another just like this one but I think i did it wrong.. here is the problem find the percent change: original-30 new- 23

OpenStudy (anonymous):

-30 and -23?

OpenStudy (anonymous):

Or 30 and 23?

OpenStudy (anonymous):

the second one

OpenStudy (anonymous):

Ok. Would you mind showing me your work?

OpenStudy (anonymous):

lol I just plugged in the numbers to the equation you gave me for the last one.

OpenStudy (anonymous):

-98.6956.....

OpenStudy (anonymous):

oh let me re-do it

OpenStudy (anonymous):

Sorry, wait a second. 23 is the new amount

OpenStudy (anonymous):

I got 30.4347......

OpenStudy (anonymous):

i rounded and got 30 but it said that was wrong.

OpenStudy (anonymous):

So sorry. Should be \(\frac{23}{30}\times 100 - 100\)

OpenStudy (anonymous):

It is negative, so it decreases

OpenStudy (anonymous):

-23.3333333333333333333333333333333333... lol

OpenStudy (anonymous):

so my percent change is -23??

OpenStudy (anonymous):

Yes

OpenStudy (anonymous):

a decrease of 23.33%

OpenStudy (anonymous):

it said that wrng

OpenStudy (anonymous):

Can you attach a screenshot?

OpenStudy (anonymous):

yes 1 sec